OpenStudy (anonymous):

solve for x: x-3/4 = 2 A) 11 B) 9 C) 5 D) 6

OpenStudy (anonymous):

xx= 11/4

OpenStudy (australopithecus):

I assume you mean 

OpenStudy (anonymous):

@Australopithecus Yes

OpenStudy (anonymous):

x-3/4=2 x=2+3/2

OpenStudy (australopithecus):

multiply both sides by 4 to eliminate the 4 from the denominator on the left side

OpenStudy (australopithecus):